### v1

# Rheologically tuned modes of collective transport in active viscoelastic films

Abstract
While many living biological media combine both viscous and elastic properties, most theoretical studies employ either purely fluid- or solid-like descriptions. We here use a unified framework for active films on substrates capable of describing a broad range of viscoelastic behavior to explore the interplay between activity and rheology. The core of the study is a comprehensive state diagram showing a rich world of spatiotemporal dynamic states. Our results demonstrate the potential of tunable rheology to realize modes of controlled active transport on the microscale.
Many living biological materials simultaneously exhibit properties of both viscous fluids and elastic solids [1] or may even change their behavior dynamically. For example, during biofilm formation, bacteria excrete extracellular polymeric substances [2, 3], which results in complex rheological behavior [4]. Depending on the stage of biofilm formation, the location within it, and the time scales considered, the material is rather fluid- or solid-like or somewhere in between. From a physical perspective, such biological systems are described within the framework of active matter [5, 6]. It concerns the collective behavior of active, self-propelled agents, frequently embedded in a surrounding medium. However, most studies refer to one of two categories, depending on the considered long-term flow behavior: active fluids or active solids.
Active fluids lack long-term elastic restoring forces and are capable of terminal flow. Primary examples are fluid suspensions of microswimmers [7, 8], such as bacteria or artificial self-propelled particles, and active nematics [9]. As described by the seminal Vicsek model [10] and its continuum counterpart, the Toner–Tu theory [11, 12], collections of active agents capable of unrestricted long-term movement may develop polar orientational order and directed collective motion. In addition to these flocking or swarming states [13, 14], active fluids exhibit various spatiotemporal patterns [15, 16, 17] including active turbulence [18, 19, 20, 21].
In contrast to active fluids, active solids are unable to support terminal flow due to elastic restoring forces. Instead, they exhibit reversible elastic deformations. In theoretical descriptions, the continuous nature of the embedding medium is often reduced to elastic spring networks of connected active agents [22, 23, 24, 25, 26]. Other examples of active solids are driven by contractile or extensile active elements and resulting stresses within the medium [27, 28], as in muscular tissue [29]. Recent topics of interest in active solids include the nature of excitations [30, 31] and the phenomenon of odd elasticity [32, 33].
There are numerous theoretical studies on either active solid-like [34, 28, 32] or active fluid-like materials with simple [12, 35, 20] or complex [36, 37, 38, 39, 40] rheologies. However, a comprehensive exploration of the spatiotemporal dynamics covering these two limiting cases and the broad spectrum in between is yet to be presented.
We aim to bridge this gap and focus on the interplay of viscous and elastic properties in active media on substrates. To this end, we employ a unified description that can be tuned continuously from the limit of active fluid-like to active solid-like properties. This framework, which covers a broad range of viscoelastic behavior in between, is derived in a recent article [41]. It is motivated from a generic microscopic picture of active agents and combined with a memory-based continuum description of the embedding medium. Basic, spatially uniform analytical solutions for resulting dynamic states were obtained for small idealized systems.
Here, we leave this idealization and present a comprehensive dynamical state diagram. It reveals fascinating and novel dynamic states when varying the activity of the suspension and the rheological properties from fluid-, via viscoelastic, to solid-like. Numerical calculations uncover complex spatiotemporal states, including stripe-like patterns in the active viscoelastic regime and separated oscillating domains for increasingly solid-like systems. Moreover, we illustrate how tuning the rheological properties in such active suspensions can be exploited for active collective transport on the microscopic scale. It allows for intermediate storage times, maintaining the spatial arrangement within a set of cargo units.
In short, in the theoretical framework, the behavior is governed by the interplay of three continuum fields [41]. The coarse-grained polar order parameter field $\mathbf{P}$ characterizes the local alignment of active units. It couples to the overall velocity field $\mathbf{v}$ and to the elastic displacement field $\mathbf{u}$. The latter captures the deformation history of the film.
First, the evolution of $\mathbf{P}$ in rescaled units follows as
$\partial_{{t}}\mathbf{P}+{\mathbf{v}}\cdot{\nabla}\mathbf{P}$ $=$ ${}-\mathbf{P}+{\nabla}^{2}\mathbf{P}+{\boldsymbol{\Omega}}\cdot\mathbf{P}$ (1)
${}+{\gamma}_{\mathrm{a}}{\mathbf{v}}\cdot[(2+\mathbf{P}\cdot\mathbf{P})\mathbf{I}/3-\mathbf{P}\mathbf{P}]\,,$
where $\mathbf{I}$ is the unit matrix. In addition to advection via the flow field $\mathbf{v}$, Eq. (1) includes reorientation via the vorticity tensor $\boldsymbol{\Omega}=[(\nabla\mathbf{v})^{\top}-(\nabla\mathbf{v})]$, where denotes the transpose. The first two terms on the right-hand side arise due to orientational and translational diffusion, while the diffusion coefficients were used to rescaled time and space. The term proportional to the alignment strength $\gamma_{\mathrm{a}}$ captures the tendency of active units to align with the overall flow field $\mathbf{v}$ near a supporting substrate [42, 43, 44, 45]. Similar terms are included in various discretized models of active solids [46, 47, 30, 48]. Nonlinearity limits the magnitude of the polar order parameter to $|\mathbf{P}|\leq 1$.
Second, the velocity field in the low-Reynolds-number limit is determined via the force balance
$\mathbf{0}=-{\nabla}{p}+{\eta}{\nabla}^{2}{\mathbf{v}}+{\mu}{\nabla}^{2}{\mathbf{u}}-{\nu}_{\mathrm{v}}{\mathbf{v}}-{\nu}_{\mathrm{d}}{\mathbf{u}}+\nu_{\mathrm{p}}\mathbf{P}\,.$ (2)
We here assume an incompressible material, $\nabla\cdot\mathbf{v}=0$, enforced by the pressure field $p$. Viscosity $\eta$ and shear modulus $\mu$ together with the displacement field $\mathbf{u}$ quantify viscous and elastic forces. Underlying friction with the substrate is set by $\nu_{\mathrm{v}}$ and slows down movement. Similarly, elastic restoring forces quantified by $\nu_{\mathrm{d}}$ result from interactions with the substrate and pull material elements back to their memorized positions. Active forcing drives the dynamics of the film and is set by $\nu_{\mathrm{p}}$.
Third, the displacement field is governed by
$\partial_{t}\mathbf{u}+\mathbf{v}\cdot\nabla\mathbf{u}=\mathbf{v}-\tau_{\mathrm{d}}^{-1}\mathbf{u}-\nabla q\,.$ (3)
While on the right-hand side overall motion $\mathbf{v}$ generates distortions, the memory of displacement $\mathbf{u}$ decays with a relaxation time $\tau_{\mathrm{d}}$. This relaxation process implies a net flow and thus incomplete elasticity. Full elastic reversibility for solids is found in the limit $\tau_{\mathrm{d}}\to\infty$. Finite values $\tau_{\mathrm{d}}$ describe a broad range of viscoelastic fluids, while the limit $\tau_{\mathrm{d}}\to 0$ refers to purely viscous fluids. Thus, rheology can be tuned continuously from fluid- to solid-like behavior. Equation (3) includes convective transport $\mathbf{v}\cdot\nabla\mathbf{u}$ by overall flow. In combination with Eq. (2), it represents a first step beyond linear elasticity [49]. This description can be derived from a generic continuum theory [50] based purely on conservation laws and symmetry arguments upon linearization in $\nabla\mathbf{u}$ [41]. The contribution $-\nabla q$ there arises from incompressibility $\nabla\cdot\mathbf{u}=0$ under linearized strains [41].
Figure 1: Dynamical state diagram as a function of the displacement relaxation time $\tau_{\mathrm{d}}$ and strength of active forcing $\nu_{\mathrm{p}}$. From left to right, the influence of elasticity increases. Colors indicate regions of different stable dynamical states. Insets in the corners illustrate idealized tracer trajectories. Blue and red dots refer to the dynamical states explored in more detail in Figs. 2 and 3, respectively. Remaining parameter values are set to $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, and $\nu_{\mathrm{d}}=10$.
Previous derivation of the theory suggested three analytical, idealized, spatially uniform mathematical solutions [41]. We here challenge their stability, amended by numerical simulations. For technical details see the Supplemental Material [51], which includes Refs. 52, 53, 54. The central goal is to construct the dynamical state diagram for active films of viscoelastic rheology.
Figure 1 shows a corresponding illustration. There, from bottom to top, we increase the strength of active forcing $\nu_{\mathrm{p}}$, starting from passive systems ($\nu_{\mathrm{p}}=0$). From left to right, with increasing displacement relaxation time $\tau_{\mathrm{d}}$, we modify the system from rather viscous, fluid-like behavior towards an increasingly elastic, solid-like system.
Indeed, on the left of Fig. 1, for $\tau_{\mathrm{d}}\rightarrow 0$ and fluid-like systems, we recover the generic stationary states. At low activity, this is a macroscopically disordered, isotropic state ($\mathbf{P}=\mathbf{0}=\mathbf{v}$). With increasing activity, it transitions to an orientationally ordered polar state ($\mathbf{P}\neq\mathbf{0}\neq\mathbf{v}$) of straight motion (see the inset on the top left). Both states in the marked regimes are linearly stable [51]. Corresponding phenomenologies are well known, for example, from the Vicsek model [10], the Toner–Tu theory [11, 12], and related extensions [19, 17, 21].
Raising the influence of elasticity by increasing $\tau_{\mathrm{d}}$, we find that both states extend to the right in Fig. 1. Yet, elasticity seems to hinder motion. Illustratively, the elastic restoring forces pin the active agents. Therefore, the phase boundary between polar moving (green) and isotropic resting (yellow) areas turns towards higher activity. The phase boundary results from linear stability analysis and is confirmed numerically [51].
The picture changes notably when we increase elasticity at elevated levels of activity from the polar state. Then novel, qualitatively different behavior emerges. It is marked by the blue region in Fig. 1. Numerical evaluations indicate a stripe-like dynamic state, see Fig. 2(a). The inset depicts the associated wiggling, snake-like motion.
Figure 2: Stripe pattern associated with the blue region in Fig. 1. (a) Snapshot of the vorticity field $\omega=(\nabla\times\mathbf{v})_{z}$ in a system of size $200\times 200$ at $\tau_{\mathrm{d}}=0.25$, see the blue dot in Fig. 1. Small dark arrows indicate the velocity field $\mathbf{v}$. Transport of the patterns, given by the phase velocity $v_{\mathrm{ph}}=-\mathrm{Im}\{\lambda(k)\}/k$, is opposite to the averaged active flow direction $\langle\mathbf{v}\rangle$. Associated material trajectories are snake-like, see the top right inset. (b) Largest growth rates $\mathrm{Re}\{\lambda(k)\}$ of modes $\mathbf{k}\|\langle\mathbf{v}\rangle$ when starting from a uniform polar state, and phase velocity $v_{\mathrm{ph}}=-\mathrm{Im}\{\lambda(k)\}/k$. Remaining parameter values are $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, $\nu_{\mathrm{d}}=10$, and $\nu_{\mathrm{p}}=20$.
Indeed, the state of globally uniform polar orientational order becomes linearly unstable in this region [51]. The growth rate $\mathrm{Re}\{\lambda(k=|\mathbf{k}|)\}$ is plotted in Fig. 2(b), with wavevectors $\mathbf{k}$ oriented parallel to $\mathbf{P}$. Importantly, the fastest-growing mode is of finite wavenumber $k_{\mathrm{max}}>0$. Thus, the instability features a certain length scale $\Lambda=2\pi/k_{\mathrm{max}}$, at least close to its onset. This length is reflected by spatial modulations along the migration direction, see Fig. 2(a). There, the local velocity field is indicated by the dark arrows. They deviate periodically in space to the left and to the right of the average migration direction given by $\langle\mathbf{v}\rangle$. The periodic modulation is further visualized when we calculate the vorticity field $\omega=(\nabla\times\mathbf{v})_{z}$, see the colored stripe pattern in Fig. 2(a).
Moreover, from $\mathrm{Im}\{\lambda(k)\}$, we calculate the phase velocity of the pattern, $v_{\mathrm{ph}}=-\mathrm{Im}(\lambda)/k$. It characterizes the traveling speed of the individual waves. Interestingly, we always find $v_{\mathrm{ph}}<0$. Consequently, the emerging structures travel in the direction opposite to global motion $\langle\mathbf{v}\rangle$, see Fig. 2(a). Supplemental Movie 1 provides a dynamic illustration.
Further increasing the influence of elasticity via $\tau_{\mathrm{d}}$ to the right in Fig. 1, we cross the red line. It marks the existence of a basic analytical solution, corresponding to rotating spatially uniform polar order $\mathbf{P}$, flow $\mathbf{v}$, and displacement $\mathbf{u}$ [51]. Yet, we do not observe this state in our spatially extended numerical systems.
Here, numerically, we identify a hysteretic regime until we reach the dotted line in Fig. 1 [51]. In this area, the emergent state depends sensitively on the initial conditions. On the one hand, upon initialization by a weakly perturbed state of global polar orientational order, still the stripe patterns develop. On the other hand, upon initialization by a weak perturbation of the above-mentioned analytical solution [51], rotations of the three fields are maintained. Further to the right of the dotted line in Fig. 1, the stripes are not found as a final dynamic state any more. Although transients can be observed for different initializations, at the end, we find a state of locally rotating fields $\mathbf{P}$, $\mathbf{v}$, and $\mathbf{u}$, see Fig. 3(a). This state is not uniform, as the system splits up into different domains of rotation. Thus, it becomes spatially heterogeneous.
Physically, the elastic restoring forces for elevated $\tau_{\mathrm{d}}$ become too strong to allow for persistent motion. The globally ordered polar or stripe-like dynamic states are associated with continuous directed transport. However, this is suppressed by higher levels of elasticity. Instead, local rotations of the polar order $\mathbf{P}$, flow field $\mathbf{v}$, and displacement field $\mathbf{u}$ combine active transport with local elastic anchoring. Imperfect circular trajectories arise, see the top right insets in Figs. 1 and 3(a).
Figure 3: Heterogeneous state of imperfect local circling motion. (a) At each spot, the vectors of polar orientational order $\mathbf{P}$ and collective flow $\mathbf{v}$ rotate over time. Dark arrows indicate $\mathbf{v}$. Domains are identified by the field of phase shift $\Delta\phi_{Pv}$ between $\mathbf{P}$ and $\mathbf{v}$, see the color bar. Positive and negative values are connected to clockwise and counterclockwise rotation, respectively. A sample trajectory is shown in the top right corner. The size of the snapshot is $100\times 100$. (b) Averaged spatial correlation function $C_{\Delta\phi_{Pv}}(\Delta x)$ of the phase shift $\Delta\phi_{Pv}$ between any two points of evaluation with distance $\Delta x$ from each other. The associated characteristic length scale $\ell$ of the dynamic patterns is identified with the first minimum of $C_{\Delta\phi_{Pv}}(\Delta x)$. Inset: Typical domain size $\ell$ as a function of active force strength $\nu_{\mathrm{p}}$ for different values of the relaxation time $\tau_{\mathrm{d}}$. Remaining parameter values are $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, $\nu_{\mathrm{d}}=10$, and $\nu_{\mathrm{p}}=20$.
The rotations of the fields $\mathbf{P}$, $\mathbf{v}$, and $\mathbf{u}$ are associated with local phase shifts between them. These phase shifts can be used to identify the local domains of rotation, see Fig. 3(a). For example, the field of phase shift $\Delta\phi_{Pv}=\phi_{P}-\phi_{v}$ refers to the local angle between $\mathbf{P}$ and $\mathbf{v}$. Remarkably, rotations in different domains may occur with opposite sense of rotation, which is indicated by the different signs in $\Delta\phi_{Pv}$. Supplemental Movie 2 provides a dynamic visualization of the dynamic state associated with the snapshot in Fig. 3(a). From there, it also becomes evident that the domains of correlated phases are constantly evolving in time.
To quantify the length scales of correlated domains, we calculate the equal-time spatial correlation function
$C_{\Delta\phi_{Pv}}(\Delta x)=\big{\langle}\Delta\phi_{Pv}(\mathbf{x},t)\,\Delta\phi_{Pv}(\mathbf{x}+\hat{\mathbf{n}}\Delta x,t)\big{\rangle}_{\hat{\mathbf{n}},\mathbf{x},t}\,.$ (4)
Here, $\Delta x$ is the distance between any two points in space and $\hat{\mathbf{n}}$ is the direction of their separation. $\langle\dots\rangle_{\hat{\mathbf{n}},\mathbf{x},t}$ denotes the average over orientations $\hat{\mathbf{n}}$, space positions $\mathbf{x}$, and time $t$. For the dynamically evolving state depicted in Fig. 3(a), we plot $C_{\Delta\phi_{Pv}}(\Delta x)$ in Fig. 3(b). The distinct length scale $\ell$ quantifying the typical domain size is identified with the location of the first minimum of $C_{\Delta\phi_{pv}}(\Delta x)$. Indeed, we find that the length scale $\ell$ of the typical size of the correlated domains is of the same order as the thickness of the stripes $\Lambda/2$ referred to in Fig. 2(a). The inset of Fig. 3(b) shows $\ell$ as a function of the strength of active forcing $\nu_{\mathrm{p}}$ for different values of the relaxation time $\tau_{\mathrm{d}}$ within the red region of Fig. 1. Mainly, we find that $\ell$ obtains a minimum, but then increases continuously with increasing activity $\nu_{\mathrm{p}}$. Thus, the structures become increasingly correlated. Approaching the transition to the isotropic state, see below, $\ell$ diverges and the structures are highly correlated in space. Overall, we find that elasticity in terms of the relaxation time $\tau_{\mathrm{d}}$ has only little influence on $\ell$ for the considered parameters.
Finally, we address the transition across the lower boundary of the red region in Fig. 1. Increasing the strength of activity $\nu_{\mathrm{p}}$ from the bottom, the isotropic state only becomes linearly unstable at the dashed black line [51]. Indeed, we observe this state numerically up to the dashed line for weakly perturbed isotropic initialization. Conversely, dynamic states of circling are found down to the red line, if the system is initialized by the weakly perturbed idealized rotational analytical solution [51]. Again, this scenario reveals a region of hysteresis. A subcritical bifurcation is identified in this area in a simplified analytical approach, as discussed separately for the idealized, spatially uniform analytical solutions [41].
As an immediate perspective, changing the rheology of active films allows to control collective transport. For example, light and external electric fields enable remote dynamic tuning of the rheology of photorheological [55, 56, 57] or electrorheological [58] materials. Exploiting such effects facilitates externally controlled transport of cargo objects by active carrier media. Persistent polar orientational order implies directed transport, while switching to the nonpolar states facilitates intermediate storage times. Elevated activity implies increased speed of transport. It requires intermediate switching to the state of imperfect local circling during rest intervals. We numerically demonstrate this scenario in Fig. 4 by periodically changing the relaxation time $\tau_{\mathrm{d}}$, see Fig. 4(a). Resulting trajectories in Fig. 4(b) show the transport of an array of cargo objects, including intermediate storage periods, while maintaining the internal positional order within the array. Suitable preparation of the substrate can guide the transport. The approach opens new routes in microfluidics powered by the activity of microswimmers such as bacteria [59, 60, 61, 62].
In summary, we uncover a rich world of dynamic states that result from the interplay between viscoelasticity and activity. We construct the overall dynamical state diagram. For example, novel phases of active matter emerge in the form of stripe patterns and heterogeneous states of imperfect circling as an immediate consequence of activity and viscoelasticity. Several features are neither observed in purely viscous active fluids nor in purely elastic active solids.
Besides the quantitative characterization of engineered active systems to control directed transport of cargo on the microscale, our description relates to the complex dynamical behavior in living biological matter. Prime examples are bacterial biofilms [2, 3, 4]. They may change their rheological characteristics dynamically and continuously from purely viscous via viscoelastic to elastic as a function of time and position within the film.
Generally, we wish to stimulate by our work experimental and theoretical research on biological and artificial active materials that exploit dynamical changes in their rheological properties to gain functionality. In this sense, future research may address the further development of adaptive living materials [63], active synthetic systems [64, 65], and actuators [66, 67]. Our results demonstrate the potential of tunable viscoelasticity to accommodate varying needs and demand during application.
Figure 4: Rheologically controlled transport of an array of cargo objects. (a) We periodically vary the relaxation time $\tau_{\mathrm{d}}$ of elastic displacements between $0.25$ and $1$. Thus, we tune the degree of elasticity. (b) Resulting trajectories for an array of four transported cargo objects. Tuning rheology allows in a controlled way to switch between persistent collective transport and intermediate storage times, roughly maintaining the spatial order within the transported cargo array. Blue and red parts in (a) and (b) correspond to each other and to the blue and red regions in the state diagram in Fig. 1. The remaining parameters are set to $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, $\nu_{\mathrm{d}}=10$, and $\nu_{\mathrm{p}}=20$. The initial edges in the cargo array are of length $1$.
Acknowledgements.
We thank the Deutsche Forschungsgemeinschaft (German Research Foundation, DFG) for support through the Research Grant no. ME 3571/5-1. Moreover, A.M.M. acknowledges support through the Heisenberg Grant no. ME 3571/4-1.